OpenStudy (anonymous):

Evaluate the expression 7x2y, when x = 5 and y = 3.

Mehek (mehek14):

\(7x^2y?\)

Mehek (mehek14):

does it look like that?

OpenStudy (anonymous):

yes @Mehek14

Mehek (mehek14):

what is \(5^2?\)

OpenStudy (anonymous):

10

Mehek (mehek14):

\(5^2=5*5\)

OpenStudy (anonymous):

my answer is 210

OpenStudy (anonymous):

but i wanna be sure its correct

Mehek (mehek14):

it's not 210

Mehek (mehek14):

\(7*5^2*3\\ 5^5=5*5=25\\7*25*3\)

OpenStudy (anonymous):

525

Mehek (mehek14):

that's the answer

OpenStudy (anonymous):

thank you @Mehek14 can you help with one more?

Mehek (mehek14):

sure

OpenStudy (anonymous):

ok go in my next tab im making a new question